i think of a number, add 1/3 of it ot half of it and the result is ten. what is the number i thought of?
Answer by kapilsinghi(68)

let the number is x
x/2 + x/3 = 10
(3x+2x)/6 =10
5x = 60
x = 60/5
x = 12
kapilsinghi123@gmail.com